# LocationReasoner: Evaluating LLMs on Real-World Site Selection Reasoning

### v1

@@ Abstract
Recent advances in large language models (LLMs), particularly those enhanced through reinforced post-training, have demonstrated impressive reasoning capabilities, as exemplified by models such as OpenAI o1 and DeepSeek-R1. However, these capabilities are predominantly benchmarked on domains like mathematical problem solving and code generation—leaving open the question of whether such reasoning skills generalize to complex, real-world scenarios. In this paper, we introduce LocationReasoner, a benchmark designed to evaluate LLMs’ reasoning abilities in the context of real-world site selection, where models must identify feasible locations by reasoning over diverse and complicated spatial, environmental, and logistical constraints. The benchmark comprises over 300 carefully crafted queries of varying difficulty levels, supported by a sandbox environment with in-house tools for constraint-based location search. Extensive evaluations reveal that state-of-the-art reasoning models offer limited improvement over their non-reasoning predecessors in real-world contexts, with even the latest OpenAI o4 model failing on 30% of site selection tasks. Moreover, agentic strategies such as ReAct and Reflexion often suffer from over-reasoning, leading to worse outcomes than direct code-generation prompting. With key limitations of LLMs in holistic and non-linear reasoning highlighted, we release LocationReasoner to foster the development of LLMs and agents capable of robust, grounded reasoning in real-world decision-making tasks. Codes and data for our benchmark are available at https://github.com/miho-koda/LocationReasoner.
@@ 1 Introduction
Large language models (LLMs) have achieved tremendous success across a broad spectrum of tasks, ranging from human-like dialogue achiam2023gpt ; team2023gemini ; liu2024deepseek ; bai2023qwen ; glm2024chatglm ; touvron2023llama to agent-based simulations park2023generative ; shanahan2023role ; wang2024survey –progress largely driven by scaling data and compute sutton2019bitter ; wei2022emergent . Recent advances extend the notion of scaling beyond training, introducing inference-time scaling techniques such as chain-of-thought (CoT) prompting wei2022chain and agentic strategies like ReAct yao2023react , which enable deeper reasoning through intermediate steps and tool use. In particular, large-scale reinforcement learning has been applied during post-training to generate high-quality reasoning traces (e.g., CoT), giving rise to a new class of reasoning models such as OpenAI o1 jaech2024openai and DeepSeek-R1 guo2025deepseek . These reasoning capabilities are critical for enabling LLMs to tackle complex, real-world tasks that often require decomposing problems, exploring alternative strategies, and interacting with external tools.
@@
Benchmarks play a crucial role in driving the progress of LLMs patterson2012better , and the rapid advancement of these models has created a growing demand for more challenging and informative evaluation tasks. While general language understanding benchmarks are widespread chiang2024chatbot ; wang2024mmlu ; hendrycks2021mmlu ; rein2024gpqa ; huang2023c , assessing reasoning capabilities remains considerably more difficult—primarily due to the challenge of verifying the correctness of LLMs’ reasoning outputs. As a result, current reasoning models are predominantly evaluated on domains with easily verifiable solutions, where verifying a solution is much simpler than generating one. This includes domains such as mathematical problem solving and programming, exemplified by benchmarks like AIME aime2024 , MATH500 math500 , and CodeForces quan2025codeelo . In contrast, few benchmarks have tackled real-world reasoning scenarios, which often involve more complicated constraints than those found in math or coding. For instance, the TravelPlanner benchmark xie2024travelplanner evaluates LLM agents on practical planning tasks but does not include recent reasoning models and lacks automated verification mechanisms, relying heavily on manual annotation. Consequently, it remains unclear whether current reasoning LLMs and agents can effectively generalize to real-world reasoning tasks. This gap highlights the urgent need for a new benchmark that targets real-world reasoning while remaining scalable and annotation-free.
@@
In this work, we introduce LocationReasoner, a benchmark designed to evaluate the real-world reasoning capabilities of LLMs through the task of site selection. Site selection is a common decision-making process in practical business settings, where the objective is to identify feasible locations based on a diverse set of criteria. This task presents a significant reasoning challenge, as it requires LLMs to perform structured location search and filtering over queries that involve multiple constraints with complex logical dependencies. For instance, selecting a site for a new restaurant may demand reasoning over temporal consumption constraints and spatial transportation constraints, with various conditions often combined through Boolean logic. To comprehensively assess LLMs’ ability to reason under such practical and multifaceted conditions, we construct the LocationReasoner benchmark, which includes over 300 site selection queries spanning a wide range of difficulty levels. The benchmark is supported by a sandbox environment equipped with offline datasets and in-house tools for constraint-based location search.
@@
LocationReasoner differs fundamentally from existing benchmarks in two key aspects. First, it focuses on practical reasoning, where LLMs must decompose complex constraints into smaller executable steps and generate codes using available tools to solve each query. Second, the benchmark features deterministic and automatically verifiable queries, enabling fully scalable evaluation without human annotation, and allowing seamless extension to much larger benchmark suites.
We summarize the key contributions of this work as follows:
  • We introduce LocationReasoner, a real-world reasoning benchmark that grounds LLMs in practical site selection tasks. The benchmark includes a suite of curated datasets, toolsets, and a sandbox environment to enable out-of-the-box evaluation of LLMs in complex constraint-based reasoning.
  • We conduct a comprehensive evaluation of four major LLM families–OpenAI jaech2024openai , Gemini team2023gemini , Claude anthropic2024claude35sonnet , and DeepSeekguo2025deepseek –covering both general-purpose and reasoning-augmented models, and two representative agentic workflows, ReAct yao2023react and Reflexion shinn2023reflexion .
@@
  • Our benchmarking results reveal that current LLMs struggle to effectively handle real-world reasoning challenges. For instance, the latest OpenAI o4 model achieves only 69.99% success on site selection queries. Furthermore, reasoning-augmented models and agentic strategies provide limited gains compared to direct prompting of general models for code generation. In-depth analysis uncovers key limitations in holistic and non-linear reasoning, pointing to critical bottlenecks that need to be addressed to improve LLM performance on practical, real-world tasks.
@@ 2 LocationReasoner
Figure 1: Query generation and execution pipeline. Queries are automatically generated using predefined constraints, ranges, thresholds, and logic composition—either through rule-based or LLM-based methods. Each query is executed via a deterministic code path using interdependent in-house tools or through an LLM agent. The results are then compared to evaluate whether the agent produces the correct filtered zones.
We center the LocationReasoner benchmark around practical site selection, a problem that is easy to understand and intuitive, but also requires multi-step reasoning across spacial, economic, and infrastructural constraints. Built on real data from Boston, MA and a fixed set of in-house tools, the benchmark offers a controlled and interpretable environment to test LLMs’ ability to reason and plan under real-world constraints.
As illustrated in Figure 1, we construct the entire query generation and execution pipeline such that it can be fully automated to support scalable evaluation. By programmatically defining constraints, threshold ranges, and logical compositions, queries can be generated in bulk through either rule-based generation or LLM-based generation. The rule-based approach is rigid and syntactically precise as predefined constraints are directly encoded into executable templates. It ensures that all conditions are logically well-formed and interpretable by deterministic code. LLM-based generation introduces linguistic diversity. The system feeds the same set of structured constraints to a language model and asks it to produce free-form queries in natural language. These queries are then routed through two execution pathways: a deterministic code-based system that applies filters using in-house tools, and an LLM agent system to be evaluated that interprets and solves the same queries. The system logs and compares the final site selections from both paths, enabling large-scale benchmarking without manual annotation or intervention.
@@ 2.1 Environment Setting
Datasets. LocationReasoner utilizes data from Boston, MA for site selection. The benchmark database is constructed by integrating multiple real-world datasets to provide a rich and realistic foundation for constraint-based planning. Specifically, we construct the database by merging data from the SafeGraph dataset [^1] https://www.safegraph.com/, which includes detailed information on Points of Interest (POIs), parking facilities, and consumer spending patterns from 2019 to 2025. To enrich the spatial and demographic context, we incorporate population data from Google Places API [^2] https://developers.google.com/maps/documentation/places/web-service/overview and transportation network data from OpenStreetMap [^3] https://www.openstreetmap.org/. This integrated dataset enables agents to reason effectively about accessibility, mobility, and demand when evaluating site suitability.
Tools. To support structured reasoning, the sandbox contains a set of in-house tools that LLM agents can call during problem solving. These tools expose curated functionalities such as filtering zones, analyzing parking availability, retrieving spending patterns across time, and evaluating transportation accessibility. We provide a detailed description of each tool and its inputs/outputs in Table 1.
@@ 2.2 Query Design
We design queries to systematically evaluate an LLM agent’s ability to reason over various constraints using in-house tools. Queries are categorized into three difficulty levels, reflecting an increasing level of complexity. Difficulty is determined by the number of constraints, diversity of tools involved, the depth of reasoning required, the linguistic complexity of the query, and the necessity to synthesize and compare heterogeneous data sources across spatial and temporal dimensions. The difficulty level enables controlled benchmarking across different kinds of reasoning scenarios, ranging from straightforward constraint filtering to complex, multi-faceted decision-making. The query variants mimic the types of natural, often nuanced questions that real users might ask when selecting a site, testing the LLM’s ability to precisely extract and interpret user requests. We provide detailed descriptions and examples of different difficulty levels as follows:
  • Easy (114 queries). Easy queries involve direct calls to a single in-house tool with a simple numerical threshold or binary filter. These queries require minimal coordination between tools.
    Example: “I want to build a retro roller rink. But the zone needs at least 3 parking lots. Where should I look?”
  • Medium (102 queries). Queries at this level introduce logical dependencies between constraints and often require coordination across multiple tools. These queries combine 2 to 3 simple constraints using Boolean operators such as AND or OR, chaining together multiple easy-level conditions. Additionally, medium queries demand mathematical reasoning and often involve relationships across structural features (e.g., neighboring zones), socio-economic indicators (e.g., spending patterns), or temporal variation (e.g., comparing data from different years).
@@
  • Hard (102 queries). Hard queries consist of typically 3-6 composite constraints derived from simple or medium-level logic, combined using AND, OR, and NOT operators. These queries demand nuanced, multi-tool reasoning, trade-off analysis, and often involve exclusionary conditions. They closely resemble real-world decision-making scenarios in site selection, where users express both strong preferences and explicit disqualifiers, requiring the agent to balance multiple objectives while avoiding infeasible options.
@@ 2.3 Reasoning Approaches
  • Direct Code Generation. We provide LLMs with relevant context of available in-house tools and structured data, prompting it to directly generate executable Python code that fulfills the query.
  • ReAct (Reasoning + Acting). The ReAct framework yao2023react is adopted to guide the agent through an iterative loop of Thought-Action-Observation (TAO). In the Thought step, the agent reflects on its current state and reasons about what to do next. In the Action step, it either calls an in-house tool or runs custom Python code. Executing custom Python code is particularly useful for performing transitional logic such as filtering results, combining intermediate outputs, or applying custom calculations. The resulting output is captured in the Observation step and stored as part of the agent’s evolving internal state. These observations are continuously fed back to the model, enabling it to reason with an evolving memory of prior tool outputs and decisions. The model is also allowed to store and reference intermediate results.
  • ReAct + Reflexion. Reflexion shinn2023reflexion extends the ReAct framework by allowing the model to reflect on failed attempts such as code errors or reaching the maximum step limit without producing a result. In these cases, the model receives feedback from its prior reasoning trace and uses it to revise its approach. This added layer of self-correction helps the model identify mistakes, adjust its logic, and improve tool usage in subsequent retries, making it especially useful for handling complex, multi-constraint queries.
@@ 2.4 Automated Verification
  • Delivery Rate. Delivery Rate measures the proportion of queries for which the LLM agent successfully returns selected zones. A process is considered completed if the agent successfully produces an output without execution failure. For agents evaluated via direct code generation, the generated code must run without syntax or runtime errors. For agents evaluated under the ReAct or Reflexion framework, the agent must avoid generating incorrect or unexecutable code, prevent infinite reasoning loops, and successfully complete the task within a strict limit of 30 reasoning steps. This metric captures basic functional reliability, as failed plans cannot be evaluated for correctness.
  • Perfect Pass Rate. The Perfect Pass Rate measures the percentage of outputs in which the set of zones returned by the agent exactly matches the ground truth derived from our objective logic. A plan is considered perfect only if it satisfies all constraints and produces no extraneous or missing zones. This strict metric ensures full compliance with the query’s requirements and serves as a high-confidence indicator of planning accuracy.
  • Precision, Recall, and F1 Score. In addition to binary success metrics, we evaluate zone-level selection accuracy using standard classification metrics. Precision measures the fraction of zones selected by the agent that are actually valid according to the ground truth. Recall measures the fraction of valid zones from the ground truth that the agent successfully identified. F1 Score is the harmonic mean of precision and recall, offering a balanced view of over- and under-selection.
3 Experiments
For the direct code generation setting, we test a range of models spanning multiple providers, including DeepSeek (V3, R1), OpenAI (GPT-4o, o4-mini), Gemini (1.5, 2.5), Claude (3, 3.5). Each model receives the same prompt structure, tool descriptions, and structured input data, and is tasked with generating executable Python code that satisfies the constraints specified in each query. Notably, we cover both general and reasoning models of each LLM family, to assess whether such advertised reasoning capabilities lead to improved performance in real-world structured, constraint-driven reasoning tasks. For the ReAct and Reflexion frameworks, we focus our evaluation on OpenAI’s GPT-4o model. This setup allows us to examine the benefits of iterative reasoning, tool chaining, and self-correction over single-shot code generation approaches.
3.1 Main Results
Table 2: Performance metrics on LocationReasoner by model and difficulty level.

### v2

@@ Abstract
Recent advances in large language models (LLMs), particularly those enhanced through reinforced post-training, have demonstrated impressive reasoning capabilities, as exemplified by models such as OpenAI o1 and DeepSeek-R1. However, these capabilities are predominantly benchmarked on domains like mathematical problem solving and code generation, leaving open the question of whether such reasoning skills generalize to complex real-world scenarios. In this paper, we introduce LocationReasoner, a benchmark designed to evaluate LLMs’ reasoning abilities in the context of real-world site selection, where models must identify feasible locations by reasoning over diverse and complicated spatial, environmental, and logistic constraints. The benchmark covers carefully crafted queries of varying difficulty levels and is supported by a sandbox environment with in-house tools for constraint-based location search. Automated verification further guarantees the scalability of the benchmark, enabling the addition of arbitrary number of queries. Extensive evaluations on real-world site selection data from Boston, New York, and Tampa reveal that state-of-the-art reasoning models offer limited improvement over their non-reasoning predecessors in real-world contexts, with even the latest OpenAI o4 model failing on 30% of site selection tasks. Moreover, agentic strategies such as ReAct and Reflexion often suffer from over-reasoning, leading to worse outcomes than direct prompting. With key limitations of LLMs in holistic and non-linear reasoning highlighted, we release LocationReasoner to foster the development of LLMs and agents capable of robust, grounded reasoning in real-world decision-making tasks. Codes and data for our benchmark are available at https://github.com/miho-koda/LocationReasoner.
@@ 1 Introduction
Large language models (LLMs) have achieved tremendous success across a broad spectrum of tasks, ranging from human-like dialogue (Achiam et al., 2023; Team et al., 2023; Liu et al., 2024; Bai et al., 2023; GLM et al., 2024; Touvron et al., 2023) to agent-based simulations (Park et al., 2023; Shanahan et al., 2023; Wang et al., 2024a)—progress largely driven by data scaling and computation (Sutton, 2019; Wei et al., 2022a). Recent advances extend the notion of scaling beyond training, introducing test-time scaling techniques such as chain-of-thought (CoT) prompting (Wei et al., 2022b) and agentic strategies such as ReAct (Yao et al., 2023b), which enable deeper reasoning through intermediate steps and tool use. In particular, large-scale reinforcement learning has been applied during post-training to generate high-quality reasoning traces (e.g., CoT), giving rise to a new class of reasoning models such as OpenAI o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025). These reasoning capabilities are critical for enabling LLMs to tackle complex real-world tasks that often require decomposing problems, exploring alternative strategies, and interacting with external tools.
@@
Benchmarks play a crucial role in driving LLM progress (Patterson, 2012), and the rapid advancement of these models has created a growing demand for more challenging and informative evaluation tasks. While general language understanding benchmarks are widespread (Chiang et al., 2024; Wang et al., 2024b; Hendrycks et al., 2021a; Rein et al., 2024; Huang et al., 2023), assessing reasoning capabilities remains considerably more difficult, primarily due to the challenge of verifying the correctness of LLMs’ reasoning outputs. As a result, current reasoning models are predominantly evaluated on domains with easily verifiable solutions, where verifying a solution is much simpler than generating one. This includes domains such as mathematical problem solving and programming, exemplified by benchmarks such as AIME (MAA, 2024), MATH500 (Hendrycks et al., 2021b), and CodeForces (Quan et al., 2025). In contrast, few benchmarks have tackled real-world reasoning scenarios, which often involve more complicated constraints. For instance, the TravelPlanner benchmark (Xie et al., 2024) evaluates LLM agents on practical planning tasks but does not include recent reasoning models and lacks automated verification mechanisms, relying heavily on manual annotation. Consequently, it remains unclear whether current reasoning LLMs and agents can effectively generalize to real-world reasoning tasks. This gap highlights the urgent need for a new benchmark that targets real-world reasoning while remaining scalable and annotation-free.
@@
In this work, we introduce LocationReasoner, a benchmark designed to evaluate the real-world reasoning capabilities of LLMs through the task of site selection. Site selection is a common decision-making process in practical business settings, where the objective is to identify feasible locations based on a diverse set of criteria. This task presents a significant reasoning challenge, as it requires LLMs to perform structured location search and filtering over queries that involve multiple constraints with complex logical dependencies. For instance, selecting a site for a new restaurant may demand reasoning over temporal consumption constraints and spatial transportation constraints, with various conditions often combined through Boolean logic. To comprehensively assess LLMs’ ability to reason under such practical and multifaceted conditions, we construct the LocationReasoner benchmark, which includes site selection queries spanning a wide range of difficulty levels. The benchmark is supported by a sandbox environment equipped with offline datasets and in-house tools for constraint-based location search.
@@
LocationReasoner differs fundamentally from existing benchmarks in two key aspects. First, it focuses on practical reasoning, where LLMs must decompose complex constraints into smaller executable steps and utilize available tools to solve each query. Second, the proposed benchmark features deterministic and automatically verifiable queries, enabling fully scalable evaluation without human annotation, and allowing seamless extension to much larger benchmark suites.
@@
The key contributions of this work are: (1) We introduce LocationReasoner, a real-world reasoning benchmark that grounds LLMs in practical site selection tasks. The benchmark includes a suite of curated datasets, toolsets, and a sandbox environment to enable out-of-the-box evaluation of LLMs in complex constraint-based reasoning. (2) We conduct a comprehensive evaluation of four major LLM families‚—OpenAI (Jaech et al., 2024), Gemini (Team et al., 2023), Claude (Anthropic, 2024), and DeepSeek (Guo et al., 2025) —covering both general-purpose and reasoning-augmented models, and two representative agentic workflows, ReAct (Yao et al., 2023b) and Reflexion (Shinn et al., 2023). (3) Our benchmarking results reveal that current LLMs struggle to effectively handle real-world reasoning challenges. For instance, the latest OpenAI o4 model achieves only 69.99% success on site selection queries. Furthermore, reasoning-augmented models and agentic strategies provide limited gains compared to direct prompting of general models. In-depth analysis uncovers key limitations in holistic and nonlinear reasoning, pointing to critical bottlenecks that need to be addressed to improve LLM performance on practical real-world tasks.
@@ 2 LocationReasoner
Figure 1: Execution pipeline.
We center the LocationReasoner benchmark around practical site selection, a problem that is easy to understand and intuitive, but also requires multi-step reasoning across all constraints. As illustrated in Figure 1, we construct the entire query generation and execution pipeline so that it can be fully automated to support scalable evaluation. Queries are generated through rule-based and LLM-based approaches to cover diverse test scenarios, using real data from Boston, New York, and Tampa. A fixed set of in-house tools are provided to achieve a controlled and interpretable environment for testing the proposed LLMs’ ability to reason and plan under real-world constraints. The queries are routed through two execution pathways: a deterministic code-based system that applies filters using in-house tools, and an LLM agent system that interprets and solves the same queries. The system logs and compares the final site selections from both paths, enabling large-scale benchmarking without manual annotation or intervention. It is worth noting that LocationReasoner supports fully automated query generation and verification, thus the benchmark can be easily scaled up to incorporate an arbitrary number of queries.
@@ 2.1 Environment Setting
Datasets. The benchmark database is constructed by integrating multiple real-world datasets to provide a rich and realistic foundation for constraint-based planning. Specifically, we construct the database by merging data from the SafeGraph dataset [^1] https://www.safegraph.com/, which includes detailed information on Points of Interest (POIs), parking facilities, and consumer spending patterns from 2019 to 2025. To enrich the spatial and demographic context, we incorporate population data from Google Places API [^2] https://developers.google.com/maps/documentation/places/web-service/overview and transportation network data from OpenStreetMap [^3] https://www.openstreetmap.org/. This integrated dataset enables agents to reason about accessibility, mobility, and demand when evaluating site suitability.
Tools. To support structured reasoning, the sandbox contains a set of in-house tools that LLM agents can call during problem solving. These tools expose curated functionalities such as filtering zones, analyzing parking availability, retrieving spending patterns over time, and evaluating transportation accessibility. We provide a detailed description of each tool in Table 1.
@@ 2.2 Query Design
We design queries to systematically evaluate an LLM agent’s ability to reason over various constraints using in-house tools. By defining constraints, threshold ranges, and logical compositions, queries can be generated in bulk through either rule-based generation or LLM-based generation. The rule-based approach is rigid and syntactically precise, as predefined constraints are directly encoded into executable templates. It ensures that all conditions are logically well-formed and interpretable by deterministic code. LLM-based generation introduces linguistic diversity. The system feeds the same set of structured constraints to a language model and asks it to produce free-form queries in natural language. Queries are categorized into three difficulty levels, reflecting an increasing level of complexity. Specifically, difficulty is determined by the number of constraints, diversity of tools involved, the depth of reasoning required, the linguistic complexity of the query, and the necessity to synthesize and compare heterogeneous data sources across spatial and temporal dimensions. The difficulty level enables controlled benchmarking across different kinds of reasoning scenarios, ranging from straightforward constraint filtering to complex, multi-faceted decision-making. The query variants mimic the types of natural, often nuanced questions that real users might ask when selecting a site, testing the LLM’s ability to precisely extract and interpret user requests. We provide detailed descriptions and examples of different difficulty levels as follows:
  • Simple. Simple queries involve direct calls to a single in-house tool with a numerical threshold or binary filter. These queries require minimal coordination between tools.
    Example: “I want to build a retro roller rink with at least 3 parking lots. Where should I look?”
  • Medium. Queries at this level introduce logical dependencies between constraints and often require coordination across multiple tools. These queries combine 2 to 3 simple constraints using Boolean operators such as AND or OR, chaining together multiple easy-level conditions. Additionally, medium queries demand mathematical reasoning and often involve relationships across structural features (e.g., neighboring zones), socio-economic indicators (e.g., spending patterns), or temporal variation (e.g., comparing data from different years).
@@
  • Hard. Hard queries consist of typically 3-6 composite constraints derived from simple or medium-level logic, combined using AND, OR, and NOT operators. These queries demand nuanced, multi-tool reasoning, trade-off analysis, and often involve exclusionary conditions. They closely resemble real-world decision-making scenarios in site selection, where users express both strong preferences and explicit disqualifiers, requiring the agent to balance multiple objectives while avoiding infeasible options.
@@ 2.3 Reasoning Approaches
  • Direct Prompting. We provide LLMs with relevant context of available in-house tools and structured data, prompting them to directly generate executable Python code that fulfills the query and outputs a list of candidate locations.
  • ReAct (Reasoning + Acting). The ReAct framework (Yao et al., 2023b) is adopted to guide the agent through an iterative loop of Thought-Action-Observation (TAO). In the Thought step, the agent reflects on its current state and reasons about what to do next. In the Action step, it either calls an in-house tool or runs custom Python code. Executing custom Python code is particularly useful for performing transitional logic such as filtering results, combining intermediate outputs, or applying custom calculations. The resulting output is captured in the Observation step and stored as part of the agent’s evolving internal state. These observations are continuously fed back to the model, enabling it to reason with an evolving memory of prior tool outputs and decisions. The model is also allowed to store and reference intermediate results.
  • ReAct + Reflexion. Reflexion (Shinn et al., 2023) extends the ReAct framework by allowing the model to reflect on failed attempts such as code errors or reaching the maximum step limit without producing a result. In these cases, the model receives feedback from its prior reasoning trace and uses it to revise its approach. This added layer of self-correction helps the model identify mistakes, adjust its logic, and improve tool usage in subsequent retries, making it especially useful for handling complex, multi-constraint queries.
@@ 2.4 Automated Verification
  • Delivery Rate measures the proportion of queries for which the LLM agent successfully returns selected zones. A process is considered completed if the agent successfully produces an output without execution failure. For agents evaluated via direct prompting, the generated code must run without syntax or runtime errors. For agents evaluated under the ReAct or Reflexion framework, the agent must avoid generating incorrect or unexecutable code, prevent infinite reasoning loops, and successfully complete the task within a strict limit of 30 reasoning steps. This metric captures basic functional reliability, as failed plans cannot be evaluated for correctness.
  • Perfect Pass Rate measures the percentage of outputs in which the set of zones returned by the agent exactly matches the ground truth derived from our objective logic. A plan is considered perfect only if it satisfies all constraints and produces no extraneous or missing zones. This strict metric ensures full compliance with the query’s requirements and serves as a high-confidence indicator of planning accuracy.
  • Precision, Recall, and F1 Score are standard classification metrics assessing zone-level selection accuracy in addition to binary success metrics.
3 Benchmarking Results
For the direct prompting setup, we test a range of models spanning multiple providers, including DeepSeek (V3, R1), OpenAI (GPT-4o, o4-mini), Gemini (1.5, 2.5), and Claude (3, 3.5). Each model receives the same prompt structure, tool descriptions, and structured input data, and is tasked with generating executable Python code that satisfies the constraints specified in each query and outputs a list of candidate locations. Notably, we cover both general and reasoning models of each LLM family, to assess whether such advertised reasoning capabilities lead to improved performance in real-world structured, constraint-driven reasoning tasks. For the ReAct and Reflexion frameworks, we focus our evaluation on OpenAI’s GPT-4o model. This setup allows us to examine the benefits of iterative reasoning, tool chaining, and self-correction over single-shot prompting approaches.
Table 2: Performance metrics on LocationReasoner by model and difficulty level on Boston data.